# Data Distribution-based Curriculum Learning

### v1

@@ III-A Density based DDCL
Figure 1(a) illustrates the DDCL (Density) process where the scoring is applied to the quantiles based on the number of samples in the quantile. Quantiles with the greatest number of samples (Q3 - 86) are given a higher score while quantiles with the least number of samples (Q1 - 12) are given a lower score. Higher scores are considered simple while lower scores are considered difficult. This results in a final quantile order sorted from the highest to lowest density thus defining the curriculum to be used during training.
[…]
2 foreach $q\in Q^{s}$ do
3       $C_{s}=C_{s}\cup|q|$
4 end foreach
[…]
return $R_{s}$
@@
Algorithm 2 provides details of the DDCL (Density) scoring process. It shows that the process starts with determining the density of each quantile by taking their cardinality, $|q|$. Once the densities for all quantiles ($C_{s}$) are known, the training data is sorted according to their quantile’s density from highest to lowest. Finally, the rearranged data, $R_{s}$, is ready for use in training.
@@ III-B Point based DDCL
On the other hand, Figure 1(b) illustrates how DDCL (Point) assigns a score to the individual data samples in the dataset by utilising their normalised Euclidean distance from centroid rather than examining quantiles. In this scoring method, data samples with the shorter Euclidean distances (P2, P5) are given the highest scores while samples with the longer Euclidean distances (P7, Pi) are assigned the lowest scores. As with DDCL (Density), higher scores in this point scoring method are considered simple while lower scores are considered difficult. Using DDCL (Point) results in a curriculum where the training samples are ordered based on their individual characteristic rather than as a group.
[…]
return $R_{s}$
@@
Algorithm 3 explains the DDCL (Point) scoring method and it shows that the data samples are sorted using the normalised Euclidean distance from centroid, $\hat{E_{s}}$. The sorting is done from the shortest distance to the longest regardless of which quantile a sample is assigned to. For example, consider a dataset with four quantiles with each quantile having at least five data samples. The first sample with the lowest $\hat{E_{s}}$ value may belong to quantile 1. Then, the sample with the second lowest $\hat{E_{s}}$ could come from quantile 4. The third lowest $\hat{E_{s}}$ value can be a sample from quantile 2. Next, the fourth lowest $\hat{E_{s}}$ may come from a sample in quantile 3. This method of sorting then repeats until $\hat{E_{s}}$ values from all data samples are accounted for with the resulting order of quantiles and samples ($R_{s}$) becoming randomised in contrast to DDCL (Density).
@@ IV Experiments
DDCL is evaluated by performing classification tasks on 7 datasets obtained from the UCI Machine Learning Repository [35]. Table I lists the datasets used for the experiments. Three types of classifiers are used for evaluating DDCL; neural network, SVM and random forest.
[b] Dataset Data Instances Attributes Class Type Breast Cancer Diagnostic 569 31 Binary Cancer 457 9 Binary Haberman 306 3 Binary Liver Disorder 346 6 Binary Pima Indians Diabetes 768 8 Binary New-Thyroid 215 5 Multi-class Diabetes 1301 86556 46 Multi-class
TABLE I: Datasets used for evaluating DDCL.
  • 1
    Preprocessed version found in [36].
The neural network architecture used for evaluating DDCL consists of one input layer, a variable number of hidden layers and one output layer. The size of the input layer is equal to the number of attributes in the dataset and varies depending on the selected dataset. Similarly, the number of hidden layers changes based on the dataset and is a hyperparameter that can be tuned. Bayesian optimisation is used on the model to tune the hidden layer values. Finally, the size of the output layer can be configured for binary or multi-class classification. The SVM classifier for evaluating DDCL uses a C value of 1.0 with a RBF kernel. The gamma ($\gamma$) value of the RBF kernel scales according to the dataset used. A RBF kernel is chosen due to its good generalisation ability and robustness to input noise. The random forest classifier in our evaluation of DDCL uses 100 estimators and a value of 2 features for splitting a node.
@@
For each dataset and classifier, 30 experiments runs are performed using three test scenarios: No Curriculum, DDCL (Density) and DDCL (Point). The first No Curriculum test scenario takes the input data and uses it to train a neural network, SVM and random forest classifier without altering the order of the training data. After training, classification experiments are conducted on the testing subset of data. In the second DDCL (Density) test scenario, the input data is processed using the steps outlined in the DDCL method and scored by density. Classification is then performed on the testing subset as stated earlier. The DDCL (Point) test scenario processes the input data through DDCL with the point scoring approach being used for the scoring step. Experiments for classification are likewise performed on the testing subset.
TABLE II: Neural Network classification results.
TABLE III: SVM classification results.
Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
Breast Cancer Diagnostic No Curriculum 91.304 100.000 96.232 $\pm$2.442
DDCL (Density) 89.130 100.000 96.594 $\pm$2.673
DDCL (Point) 86.957 100.000 97.319 $\pm$2.953
Cancer No Curriculum 86.486 100.000 94.955 $\pm$3.993
DDCL (Density) 91.667 100.000 96.852 $\pm$2.658
DDCL (Point) 86.111 100.000 96.574 $\pm$3.100
Haberman No Curriculum 45.833 79.167 65.833 $\pm$8.898
DDCL (Density) 50.000 87.500 69.167 $\pm$10.353
DDCL (Point) 41.667 87.500 67.222 $\pm$9.424
Liver Disorder No Curriculum 57.143 85.714 69.048 $\pm$6.477
DDCL (Density) 46.429 85.714 68.214 $\pm$9.052
DDCL (Point) 57.143 82.143 69.524 $\pm$6.098
Pima Indians Diabetes No Curriculum 54.098 86.885 71.803 $\pm$6.658
DDCL (Density) 55.738 85.246 72.732 $\pm$5.933
DDCL (Point) 57.377 80.328 71.803 $\pm$5.460
New-Thyroid No Curriculum 82.353 100.000 93.137 $\pm$5.280
DDCL (Density) 82.353 100.000 95.490 $\pm$5.409
DDCL (Point) 82.353 100.000 94.118 $\pm$5.683
Diabetes 130 No Curriculum 53.235 54.694 54.029 $\pm$0.518
DDCL (Density) 56.148 56.798 56.579 $\pm$0.246
DDCL (Point) 54.400 56.278 55.139 $\pm$0.696
Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
Breast Cancer Diagnostic No Curriculum 94.737 99.415 97.700 $\pm$1.188
DDCL (Density) 95.322 99.415 97.758 $\pm$1.025
DDCL (Point) 94.737 99.415 97.719 $\pm$1.239
Cancer No Curriculum 94.928 99.275 96.715 $\pm$1.134
DDCL (Density) 94.118 99.265 96.985 $\pm$1.145
DDCL (Point) 92.647 99.265 96.225 $\pm$1.562
Haberman No Curriculum 65.217 78.261 72.609 $\pm$4.051
DDCL (Density) 65.217 83.696 73.841 $\pm$4.836
DDCL (Point) 67.391 79.348 73.080 $\pm$3.329
Liver Disorder No Curriculum 44.231 76.923 63.750 $\pm$7.264
DDCL (Density) 49.038 75.000 66.699 $\pm$5.515
DDCL (Point) 55.769 75.962 68.558 $\pm$4.783
Pima Indians Diabetes No Curriculum 71.429 80.519 76.046 $\pm$2.500
DDCL (Density) 70.996 82.251 77.128 $\pm$2.991
DDCL (Point) 72.294 82.684 76.508 $\pm$2.514
New-Thyroid No Curriculum 92.308 100.000 97.128 $\pm$2.472
DDCL (Density) 92.308 100.000 97.282 $\pm$1.892
DDCL (Point) 92.308 100.000 96.872 $\pm$1.563
Diabetes 130 No Curriculum 64.285 65.487 64.948 $\pm$0.302
DDCL (Density) 64.220 65.457 64.977 $\pm$0.298
DDCL (Point) 64.301 65.469 65.052 $\pm$0.266
TABLE III: SVM classification results.
TABLE IV: Random Forest classification results.
Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
Breast Cancer Diagnostic No Curriculum 94.152 98.246 95.945 $\pm$1.034
DDCL (Density) 92.398 98.246 95.653 $\pm$1.581
DDCL (Point) 91.228 98.246 95.945 $\pm$1.517
Cancer No Curriculum 92.754 100.000 97.150 $\pm$1.374
DDCL (Density) 94.853 100.000 96.691 $\pm$1.298
DDCL (Point) 94.118 100.000 97.206 $\pm$1.361
Haberman No Curriculum 58.696 76.087 68.261 $\pm$4.100
DDCL (Density) 65.217 79.348 70.145 $\pm$3.825
DDCL (Point) 57.609 77.174 69.746 $\pm$4.367
Liver Disorder No Curriculum 59.615 78.846 70.417 $\pm$5.104
DDCL (Density) 67.308 77.885 72.917 $\pm$2.709
DDCL (Point) 60.577 78.846 70.064 $\pm$4.077
Pima Indians Diabetes No Curriculum 71.861 79.654 75.483 $\pm$1.985
DDCL (Density) 70.996 80.087 75.339 $\pm$2.200
DDCL (Point) 65.368 81.818 76.003 $\pm$3.461
New-Thyroid No Curriculum 95.385 100.000 97.949 $\pm$1.214
DDCL (Density) 92.308 100.000 98.000 $\pm$1.911
DDCL (Point) 95.385 100.000 97.897 $\pm$1.512
Diabetes 130 No Curriculum 65.830 66.658 66.218 $\pm$0.228
DDCL (Density) 65.626 66.948 66.242 $\pm$0.309
DDCL (Point) 65.542 66.805 66.199 $\pm$0.293
The results of all three test scenarios using neural networks, SVM and random forest classifiers are given in Table III, Table III and Table IV respectively.
Figure 3: Precision-Recall curves for binary classification datasets and confusion matrix for multi-class datasets.
@@ V Discussion
Observations on the learning methods show that the classification performance across the datasets is always increased whenever DDCL is applied. This is true regardless of the learning method used as shown in Table III, Table III and Table IV.
Figure 4: Error loss per epoch for neural network classifier.
Using the neural network classifier provides the highest best results for the Haberman, Liver Disorder and Pima Indians Diabetes datasets while the SVM method achieves the highest average results on the Breast Cancer Diagnostic, Haberman and Pima Indians Diabetes data. The random forest based classifier yields the highest average result for the Cancer, Liver Disorder, New-Thyroid and Diabetes 130 datasets. Notably, using DDCL with SVM and random forest approaches results in significant increases in average accuracy for Haberman and Liver Disorder data respectively compared to other learning methods. Figure 3 shows the precision-recall curves for the best result of each binary classification dataset when using the neural network method. For multi-class data, the confusion matrix is provided in place of the precision-recall curves.
The highest average result for Breast Cancer Diagnostic was achieved using SVM with density-based DDCL while the largest improvement (1.087%) was using the neural network with point-based DDCL. For the Cancer dataset, using random forest with point-based DDCL provided the highest average accuracy and using the neural network with density-based DDCL gave the most improved accuracy (1.897%). Using SVM with density-based DDCL resulted in the highest average accuracy when evaluated on the Haberman data. The largest improvement for this dataset (3.334%) was observed when using density-based DDCL with the neural network. The highest average result for Liver Disorder was achieved using random forest with density-based DDCL while the SVM gave the largest average accuracy improvement at 4.808%. For the Pima Indians Diabetes dataset, the highest average accuracy was obtained using SVM with density-based DDCL which was also the most improved accuracy (1.082%).
In multi-class classification, using random forest with density-based DDCL resulted in the highest average accuracy on the New-Thyroid data and the largest improvement for this dataset came from density-based DDCL applied to the neural network. For the Diabetes 130 data, density-based DDCL with random forest provided the highest average accuracy while the largest improvement in accuracy (2.550%) was obtained through the neural network classifier utilising density-based DDCL.
The experiment results demonstrate that using DDCL to order the training data for a given learning method leads to improvements in classification accuracy. This can be seen in the performance of all datasets tested with either the density or point-based DDCL approach having increased performance compared to the no curriculum approach. Figure 4 shows the error loss per epoch plots when using the neural network classifier for each dataset except for Diabetes 130 which is excluded since its average accuracy is not significant [36]. The plots show that using DDCL leads to better reduction in the error loss.
In addition, using DDCL to order the training data results in faster convergence. We arrived at this conclusion by examining the error loss for the first five epochs of each dataset using batch gradient descent. Figure 5 shows the error loss plots for each dataset. It can be seen that the overall losses for DDCL (Density) and DDCL (Point) are reduced faster compared to the no curriculum approach thus leading to faster convergence towards the minimum. This is consistent with the findings of the error loss against epoch plots in Figure 4 where the error loss towards the end is lower for DDCL (Density) and DDCL (Point).

### v2

@@ III-A Density based DDCL
Figure 2(a) illustrates the DDCL-Density process where the scoring is applied to the quantiles based on the number of samples in the quantile. Quantiles with the greatest number of samples (Q6) are given a higher score whereas quantiles with the least number of samples (Q1) are given a lower score. Higher scores are considered simple and lower scores are considered difficult. This results in a final quantile order sorted from the highest to lowest density thus defining the curriculum to be used during training.
TABLE II: Neural Network results with and without DDCL.
Dataset Test Scenario Worst % Best % Average ±$\pm\sigma$
Breast Cancer (Diagnostic) No Curriculum 91.304 100.000 96.232 $\pm$2.442
DDCL-Density 89.130 100.000 96.594 $\pm$2.673
DDCL-Point 86.957 100.000 97.319 $\pm$2.953
Cancer No Curriculum 86.486 100.000 94.955 $\pm$3.993
DDCL-Density 91.667 100.000 96.852 $\pm$2.658
DDCL-Point 86.111 100.000 96.574 $\pm$3.100
Haberman’s Survival No Curriculum 45.833 79.167 65.833 $\pm$8.898
DDCL-Density 50.000 87.500 69.167 $\pm$10.353
DDCL-Point 41.667 87.500 67.222 $\pm$9.424
Liver Disorder No Curriculum 57.143 85.714 69.048 $\pm$6.477
DDCL-Density 46.429 85.714 68.214 $\pm$9.052
DDCL-Point 57.143 82.143 69.524 $\pm$6.098
Pima Indians Diabetes No Curriculum 54.098 86.885 71.803 $\pm$6.658
DDCL-Density 55.738 85.246 72.732 $\pm$5.933
DDCL-Point 57.377 80.328 71.803 $\pm$5.460
New-Thyroid No Curriculum 82.353 100.000 93.137 $\pm$5.280
DDCL-Density 82.353 100.000 95.490 $\pm$5.409
DDCL-Point 82.353 100.000 94.118 $\pm$5.683
Diabetes 130 No Curriculum 53.235 54.694 54.029 $\pm$0.518
DDCL-Density 56.148 56.798 56.579 $\pm$0.246
DDCL-Point 54.400 56.278 55.139 $\pm$0.696
TABLE III: SVM results with and without DDCL.
Dataset Test Scenario Worst % Best % Average ±$\pm\sigma$
Breast Cancer (Diagnostic) No Curriculum 94.737 99.415 97.700 $\pm$1.188
DDCL-Density 95.322 99.415 97.758 $\pm$1.025
DDCL-Point 94.737 99.415 97.719 $\pm$1.239
Cancer No Curriculum 94.928 99.275 96.715 $\pm$1.134
DDCL-Density 94.118 99.265 96.985 $\pm$1.145
DDCL-Point 92.647 99.265 96.225 $\pm$1.562
Haberman’s Survival No Curriculum 65.217 78.261 72.609 $\pm$4.051
DDCL-Density 65.217 83.696 73.841 $\pm$4.836
DDCL-Point 67.391 79.348 73.080 $\pm$3.329
Liver Disorder No Curriculum 44.231 76.923 63.750 $\pm$7.264
DDCL-Density 49.038 75.000 66.699 $\pm$5.515
DDCL-Point 55.769 75.962 68.558 $\pm$4.783
Pima Indians Diabetes No Curriculum 71.429 80.519 76.046 $\pm$2.500
DDCL-Density 70.996 82.251 77.128 $\pm$2.991
DDCL-Point 72.294 82.684 76.508 $\pm$2.514
New-Thyroid No Curriculum 92.308 100.000 97.128 $\pm$2.472
DDCL-Density 92.308 100.000 97.282 $\pm$1.892
DDCL-Point 92.308 100.000 96.872 $\pm$1.563
Diabetes 130 No Curriculum 64.285 65.487 64.948 $\pm$0.302
DDCL-Density 64.220 65.457 64.977 $\pm$0.298
DDCL-Point 64.301 65.469 65.052 $\pm$0.266
TABLE IV: Random Forest results with and without DDCL.
Dataset Test Scenario Worst % Best % Average ±$\pm\sigma$
Breast Cancer (Diagnostic) No Curriculum 94.152 98.246 95.945 $\pm$1.034
DDCL-Density 92.398 98.246 95.653 $\pm$1.581
DDCL-Point 91.228 98.246 95.945 $\pm$1.517
Cancer No Curriculum 92.754 100.000 97.150 $\pm$1.374
DDCL-Density 94.853 100.000 96.691 $\pm$1.298
DDCL-Point 94.118 100.000 97.206 $\pm$1.361
Haberman’s Survival No Curriculum 58.696 76.087 68.261 $\pm$4.100
DDCL-Density 65.217 79.348 70.145 $\pm$3.825
DDCL-Point 57.609 77.174 69.746 $\pm$4.367
Liver Disorder No Curriculum 59.615 78.846 70.417 $\pm$5.104
DDCL-Density 67.308 77.885 72.917 $\pm$2.709
DDCL-Point 60.577 78.846 70.064 $\pm$4.077
Pima Indians Diabetes No Curriculum 71.861 79.654 75.483 $\pm$1.985
DDCL-Density 70.996 80.087 75.339 $\pm$2.200
DDCL-Point 65.368 81.818 76.003 $\pm$3.461
New-Thyroid No Curriculum 95.385 100.000 97.949 $\pm$1.214
DDCL-Density 92.308 100.000 98.000 $\pm$1.911
DDCL-Point 95.385 100.000 97.897 $\pm$1.512
Diabetes 130 No Curriculum 65.830 66.658 66.218 $\pm$0.228
DDCL-Density 65.626 66.948 66.242 $\pm$0.309
DDCL-Point 65.542 66.805 66.199 $\pm$0.293
@@
1:  $C_{s}\leftarrow\varnothing$Initialize}
2:  for all $q\in Q^{s}$ do
3:     $C_{s}\leftarrow C_{s}\cup|q|$
4:  end for
5:  $R_{s}\leftarrow sort(\forall\>C_{s}\mid s\in S)$
6:
7:  return  $R_{s}$
Algorithm 2 provides details of the DDCL-Density scoring process. It shows that the process starts with determining the density of each quantile by taking their cardinality, $|q|$. Once the densities for all quantiles ($C_{s}$) are known, the training data is sorted according to their quantile’s density from highest to lowest. Finally, the rearranged data, $R_{s}$, is ready for use in training.
@@ III-B Point based DDCL
On the other hand, Figure 2(b) illustrates how DDCL-Point assigns a score to the individual data samples in the dataset by utilizing their normalized Euclidean distance from centroid rather than examining quantiles. In this scoring method, data samples with the shorter Euclidean distances (1 and 5) are given the highest scores whereas samples with the longer Euclidean distances (3 and $s$) are assigned the lowest scores. As with DDCL-Density, higher scores in this point scoring method are considered simple whereas lower scores are considered difficult. Using DDCL-Point results in a curriculum where the training samples are ordered based on their individual characteristic rather than as a group.
@@
1:  $R_{s}\leftarrow sort(\forall\>\hat{E_{s}}|s\in S)$
2:
3:  return  $R_{s}$
Algorithm 3 explains the DDCL-Point scoring method and it shows that the data samples are sorted using the normalized Euclidean distance from centroid, $\hat{E_{s}}$. The sorting is done from the shortest distance to the longest regardless of which quantile a sample is assigned to. For example, consider a dataset with four quantiles with each quantile having at least five data samples. The first sample with the lowest $\hat{E_{s}}$ value may belong to quantile 1. Then, the sample with the second lowest $\hat{E_{s}}$ could come from quantile 4. The third lowest $\hat{E_{s}}$ value can be a sample from quantile 2. Next, the fourth lowest $\hat{E_{s}}$ may come from a sample in quantile 3. This method of sorting then repeats until $\hat{E_{s}}$ values from all data samples are accounted for with the resulting order of quantiles and samples ($R_{s}$) becoming randomized in contrast to DDCL-Density.
Figure 3: Precision-Recall curves for binary classification datasets.
Figure 4: Confusion matrix for multi-class classification datasets.
@@ IV Experiments
DDCL is evaluated by performing classification tasks on seven datasets obtained from the UCI Machine Learning Repository [43]. As listed in Table I, the datasets used for the experiments focus on small to medium-sized medical data. Three types of classifiers are used for evaluating DDCL; neural network, SVM and random forest.
TABLE V: Comparison of DDCL results with other works.
Dataset Method Year Accuracy %
Breast Cancer Wisconsin (Diagnostic) SVM [44] 2021 97.20
SVM (DDCL) 2024 99.42
Haberman’s Survival Random Forest [45] 2023 74.00
Random Forest (DDCL) 2024 79.35
Liver Disorder Neural Network [46] 2023 76.67
Neural Network (DDCL) 2024 82.14
Pima Indians Diabetes Random Forest [47] 2022 79.57
Random Forest [45] 2023 79.00
Random Forest (DDCL) 2024 81.82
Diabetes 130 Random Forest [42] 2018 55.97
Random Forest (DDCL) 2024 66.95
The neural network architecture used for evaluating DDCL consists of one input layer, a variable number of hidden layers and one output layer. The size of the input layer is equal to the number of attributes in the dataset and varies depending on the selected dataset. Similarly, the number of hidden layers changes based on the dataset and is a hyperparameter that can be tuned. Bayesian optimization is used on the model to tune the hidden layer values. Finally, the size of the output layer can be configured for binary or multi-class classification. The SVM classifier for evaluating DDCL uses a C value of 1.0 with a RBF kernel. The gamma ($\gamma$) value of the RBF kernel scales according to the dataset used. A RBF kernel is chosen due to its good generalization ability and robustness to input noise. The random forest classifier in our evaluation of DDCL uses 100 estimators and a value of 2 features for splitting a node.
@@
For each dataset, 30 experiments runs are performed for each of the three test scenarios: No Curriculum, DDCL-Density and DDCL-Point. The same number of experiment runs is used for all three classifiers. The first No Curriculum test scenario takes the input data and uses it to train a neural network, SVM and random forest classifier without altering the order of the training data. After training, classification experiments are conducted on the testing subset of data. In the second DDCL-Density test scenario, the input data is processed using the steps outlined in the DDCL method and scored by density. Classification is then performed on the testing subset as stated earlier. The DDCL-Point test scenario processes the input data through DDCL with the point scoring approach being used for the scoring step. Experiments for classification are likewise performed on the testing subset.
Figure 5: Error loss per epoch for neural network classifier.
The results for neural networks, SVM and random forest classifiers are given in Table II, Table III and Table IV respectively.
@@ V Discussion
Observations on the learning methods show that the classification performance across the datasets is always increased whenever DDCL is applied. This is true regardless of the learning method used as shown in Table II, Table III and Table IV.
@@
Using the neural network classifier provides the highest best results for the Haberman’s Survival, Liver Disorder and Pima Indians Diabetes datasets whereas the SVM method achieves the highest average results on the Breast Cancer (Diagnostic), Haberman’s Survival and Pima Indians Diabetes data. The random forest based classifier yields the highest average result for the Cancer, Liver Disorder, New-Thyroid and Diabetes 130 datasets. Notably, using DDCL with SVM and random forest approaches results in significant increases in average accuracy for Haberman’s Survival and Liver Disorder data respectively compared to other learning methods. Figure 3 shows the precision-recall curves for the best result of each binary classification dataset when using the neural network method. For multi-class data, the confusion matrix is provided in place of the precision-recall curves by Figure 4.
Table V shows a comparison of results obtained using DDCL compared to approaches used by other state-of-the-art works. It demonstrates DDCL’s ability to achieve superior results compared to the same type of classifier trained using a different algorithm. In particular it outperforms the balanced stratified method used in [45] on the challenging Haberman’s Survival and Pima Indians Diabetes datasets. Furthermore, DDCL provides greater generalization capability and performance with neural networks on the Liver Disorder data compared to the standard approach used in [46].
The highest average result for Breast Cancer (Diagnostic) was achieved using SVM with density-based DDCL while the largest improvement (1.087%) was using the neural network with point-based DDCL. For the Cancer dataset, using random forest with point-based DDCL provided the highest average accuracy and using the neural network with density-based DDCL gave the most improved accuracy (1.897%). Using SVM with density-based DDCL resulted in the highest average accuracy when evaluated on the Haberman’s Survival data. The largest improvement for this dataset (3.334%) was observed when using density-based DDCL with the neural network. The highest average result for Liver Disorder was achieved using random forest with density-based DDCL whereas the SVM gave the largest average accuracy improvement at 4.808%. For the Pima Indians Diabetes dataset, the highest average accuracy was obtained using SVM with density-based DDCL which was also the most improved accuracy (1.082%).
In multi-class classification, using random forest with density-based DDCL resulted in the highest average accuracy on the New-Thyroid data and the largest improvement for this dataset (2.353%) came from density-based DDCL applied to the neural network. For the Diabetes 130 data, density-based DDCL with random forest provided the highest average accuracy while the largest improvement in accuracy (2.550%) was obtained through the neural network classifier utilizing density-based DDCL.
The experiment results demonstrate that using DDCL to order the training data for a given learning method leads to improvements in classification accuracy. This can be seen in the performance of all datasets tested with either the density or point-based DDCL approach having increased performance compared to the no curriculum approach. Figure 5 shows the error loss per epoch plots when using the neural network classifier for each dataset except for Diabetes 130 which is excluded since its average accuracy is not significant [42]. The plot for Breast Cancer (Diagnostic) shows a minimal change in the DDCL error loss trends compared to the no curriculum method whereas the Cancer dataset plot presents a reduction in the error loss when using DDCL. For the remaining datasets, the use of DDCL shows a significant and quick reduction in the error loss.
In addition, using DDCL to order the training data results in faster convergence. We arrived at this conclusion by examining the error loss for the first five epochs of each dataset using batch gradient descent. Figure 6 shows the error loss plots for each dataset. It can be seen that the overall losses for DDCL-Density and DDCL-Point are reduced faster compared to the no curriculum approach thus leading to faster convergence towards the minimum. This is consistent with the findings of the error loss against epoch plots in Figure 5 where the error loss towards the end is lower for DDCL-Density and DDCL-Point.